# Cardinality-Preserving Attention Channels for Graph Transformers in Molecular Property Prediction

### v3

@@ 6 Limitations
$K=3$ may miss long-range effects in very large molecules. Gains shown on public benchmarks; prospective validation remains future work. Our mask-based sparsity does not exploit dedicated block-sparse kernels, limiting efficiency gains to effective sequence length reduction. Near-duplicate risks persist despite scaffold decontamination. Per-view SPD increases pretraining cost.
@@ 7 Conclusion
CardinalGraphFormer advances graph transformers with a query-conditioned cardinality-preserving channel, yielding consistent gains across diverse molecular benchmarks.
@@
Code and artifacts will be released upon acceptance (reserved DOI: 10.5281/zenodo.18622116). Includes environment.yml, dependency versions, filtering/split scripts, and reproduction commands for main tables.

### v4

@@ 7 Limitations
$K=3$ may miss long-range effects in very large molecules. Gains demonstrated on public benchmarks; prospective validation on proprietary datasets remains future work. Our mask-based sparsity does not exploit dedicated block-sparse kernels, limiting efficiency gains to effective sequence length reduction.
Lipinski Filtering Bias. The Lipinski-style filtering applied during corpus construction (MW $\leq 500$, logP $\leq 5$, HBD $\leq 5$, HBA $\leq 10$) biases the pretraining distribution toward traditional drug-like chemical space. This excludes beyond-Rule-of-Five compounds including macrocycles, PROTACs, and other non-traditional chemotypes that are increasingly central to modern drug discovery. Transfer to such chemical classes may require additional fine-tuning or corpus augmentation.
Query-Only Gating. The CPA mechanism employs query-only gating ($g_{i}=\sigma(W_{g}q_{i})$) rather than per-neighbor gating ($g_{ij}=\sigma(W_{g}[q_{i};k_{j}])$). While query-only gating is computationally efficient and stable, per-neighbor gating could provide greater expressive power by modulating the unnormalized sum on a per-edge basis. This variant was not explored due to computational constraints and potential numerical instability, representing a limitation in the mechanism’s flexibility.
Near-duplicate risks persist despite scaffold decontamination. Per-view SPD computation increases pretraining cost by  15%, and the absolute cost remains non-trivial for broader adoption.
@@ 8 Conclusion
CardinalGraphFormer advances graph transformers with a query-conditioned cardinality-preserving channel grounded in expressivity principles, yielding consistent gains across diverse molecular benchmarks under matched experimental protocols.
@@
Code and artifacts are provided in the accompanying reproducibility package (environment specification, dependency versions, preprocessing/split scripts, per-seed result files, and reproduction commands for all main tables). A reserved Zenodo DOI for the archival package is 10.5281/zenodo.18622116; if required by venue policy, public access to the DOI landing page will be activated upon acceptance.
@@ A.5 Size-Shift Split Construction
Appendix B Statistical Details
Task $\Delta$ (vs. SparseGraphormer-K3) 95% Bootstrap CI Holm-adjusted p-value
ESOL (RMSE $\downarrow$) -0.056 [-0.072, -0.040] 0.002
Lipophilicity (RMSE $\downarrow$) -0.029 [-0.041, -0.017] 0.015
BBBP (ROC-AUC $\uparrow$) +0.020 [0.012, 0.028] 0.008
Tox21 (ROC-AUC $\uparrow$) +0.013 [0.007, 0.019] 0.032
ClinTox (ROC-AUC $\uparrow$) +0.015 [0.005, 0.025] 0.045
ogbg-molhiv (ROC-AUC $\uparrow$) +0.017 [0.009, 0.025] 0.012
ogbg-molpcba (AP $\uparrow$) +0.010 [0.006, 0.014] 0.028
Caco2_Wang (MAE $\downarrow$) -0.024 [-0.032, -0.016] 0.005
hERG (ROC-AUC $\uparrow$) +0.028 [0.022, 0.034] 0.001
Clearance_Hepatocyte_AZ (Spearman $\uparrow$) +0.013 [0.003, 0.023] 0.040
Clearance_Microsome_AZ (Spearman $\uparrow$) +0.007 [0.001, 0.013] 0.048
Table 16: Detailed statistics for primary deltas (5 seeds). CIs from paired bootstrap (10,000 resamples); Holm correction across 11 tasks. All adjusted p <0.05<0.05, confirming significance.